OpenStudy (anonymous):

Let f(x) = 2x+1 and g(x) = x-1 / 2. Find: f(g(3))

OpenStudy (anonymous):

Basically its f(g(3))= 2(x-1/2)+1 So its 2(3-1/2) + 1 2(6/2-1/2) + 1 2(5/2)+1 10/2 + 1 6 should be the answer try it out. This is my first time answering a question on OpenStudy hopefully this is correct.

OpenStudy (anonymous):

Yes dude. You are correct. And OpenStudy is good. Should take advantage of it. You don't get ASAP answers .... but you eventually get your answers. Sometimes people explain it to you ... then other times you get direct answers. Pretty much its good for what is taken advantage for. Hope you have fun. Welcome to OpenStudy. And thanks for your answer by the way. @arshia93

OpenStudy (anonymous):

no problem man, give me that best response so it doesnt look like all i do is ask questions on here lol @rbeckford134

OpenStudy (anonymous):

@arshia93 Dude, I have another one: f(g(3)) = 2(x-1/2) + 1 = 2(3-1/2) + 1 =2(6 - 1 /2 ) + 1 =2 ( 5 /2 ) + 1 = 10/2 + 1 = 5 + 1 = 6

OpenStudy (anonymous):

Did I complete this correctly?

OpenStudy (anonymous):

isn't this the same exact problem?

OpenStudy (anonymous):

In a way .... yes, just the opposite .... g ( f(3)) .. I did it your way so i wanted to see if my answer was correct. And if I made a mistake ... I would try to solve to make it correct.

OpenStudy (anonymous):

so now basically you plug f(3) into g(x). Keep in mind that f(3) means where there is x in f(x) plug in 3, just like how we plugged it in with g(3). Try it now Let f(x) = 2x+1 and g(x) = x-1 / 2. Find: g(f(3)) (2(3)+1)-1/2 now do the next step...

OpenStudy (anonymous):

\[\frac{ (6+1 )}{ 2 } + 1\] \[\frac{ 7 }{ 2 } + 1\] \[\frac{ 7 + 1 }{ 2 }\] \[\frac{ 8 }{ 2 }\] = 4